OpenStudy (anonymous):

|3/5x+6|-1/2=8 PLEASE HELP KEEP GETTING THE WRONG ANSWER

OpenStudy (anonymous):

try solving: 3/5*x+6 = 8 1/2 and 6-3/5*x = 8 1/2

OpenStudy (anonymous):

When you get the absolute value by itself on one side (like you would a variable), solve for the negative and positive possibilities of the variable inside of the absolute value sign. You should get more than one answer.

OpenStudy (anonymous):

the answers i got keep coming out wrong on my online homework

OpenStudy (anonymous):

what answers did you get?

OpenStudy (anonymous):

-24.16 & 4.16

OpenStudy (anonymous):

Before you split the absolute value up into 2 equations, you have to isolate it from all of the other terms, in this case you have to get rid of the 1/2 that the absolute value is being subtracted by. Then, once you have only the absolute value equaling something, plug in an x where the x is in the absolute value, take away the absolute value function and solve that equation, then plug in a -x where x is in the absolute value, take away the absolute value function, and solve that one as well. Don't forget to put parenthesis around the term you're substituting [for example, if I substitute 3x-2=7 where x=-9, I would then say 3*(-9)-2=7].